# Spectrum and Resonances of Stark Hamiltonians with $\delta$–Shell Potentials

### v1

@@ 5.4 Small–field asymptotics in three dimensions
Step 1: Setup at $F=0$. Suppose $E_{0}<0$ is a simple eigenvalue of $H_{0,\alpha}$ with boundary eigenvector $\phi_{0}^{\flat}\in L^{2}(S_{a})$ normalized by
$\bigl{(}(M_{0}^{\flat})^{\prime}(E_{0})\phi_{0}^{\flat},\phi_{0}^{\flat}\bigr{)}_{L^{2}(S_{a})}=1,\quad M_{0}^{\flat}(z)=\Lambda^{-1}\tau R_{0}(z)\tau^{*}\Lambda^{-1},\quad R_{0}(z)=(-\Delta-z)^{-1}.$ (5.3)
Step 2: Analytic expansion of $M_{F}(z)$. For $|F|$ sufficiently small, the resolvent admits the Neumann expansion
$R_{F,0}(z)=(H_{0}+Fx_{1}-z)^{-1}=R_{0}(z)-FR_{0}(z)\,x_{1}\,R_{0}(z)+F^{2}R_{0}(z)\,x_{1}R_{0}(z)\,x_{1}R_{0}(z)$-$\cdots$
valid for $z\in\mathbb{C}\setminus[0,\infty)$. Substituting into $M_{F}(z)=\tau R_{F,0}(z)\tau^{*}$ gives
$M_{F}(z)$ $=$ $M_{0}(z)+FM_{1}(z)+F^{2}M_{2}(z)+O(F^{3}),$
$M_{1}(z)$ $=$ $-\tau R_{0}(z)\,x_{1}\,R_{0}(z)\tau^{*},$
$M_{2}(z)$ $=$ $\tau R_{0}(z)\,x_{1}R_{0}(z)\,x_{1}R_{0}(z)\tau^{*}.$
Step 3: Analytic perturbation of the eigenvalue. Let $G(F,z)=I+A_{\alpha}M_{F}(z)$. Since $E_{0}$ is a simple zero of $\det(\Lambda^{-1}G(0,z)\Lambda)$, the implicit function theorem yields an analytic branch $z(F)=E_{0}+Fa_{1}+F^{2}a_{2}+O(F^{3})$ with
$a_{1}=-\bigl{(}M_{1}^{\flat}(E_{0})\phi_{0}^{\flat},\phi_{0}^{\flat}\bigr{)}_{L^{2}(S_{a})}.$ (5.4)
Step 4: Exponentially small resonance widths. Let $\phi_{0}^{\flat}\in L^{2}(S_{a})$ be the normalized boundary eigenvector corresponding to the simple eigenvalue $E_{0}<0$ of $H_{0,\alpha}$ as in Step 1. Then the branch $z(F)$ obtained in Step 3 satisfies $\text{{\rm Re}}z(F)\to E_{0}$ as $F\downarrow 0$. It is known from semiclassical tunneling theory that the imaginary part $\text{{\rm Im}}z(F)$ is exponentially small in $1/F$. To make this more explicit, we recall that the exponential factor can be computed in terms of an Agmon–type action integral along the field direction.
Agmon distance along the field (derivation).
In the exterior region $\{x_{1}>a\}$ the effective potential is $V(x_{1})=Fx_{1}$, and for the bound state energy $E_{0}<0$ (at $F=0$) the Agmon action along the straight path in the field direction is
$d(a\to x_{1})=\int_{a}^{x_{1}}\!\sqrt{Fs-E_{0}}\,ds=\frac{2}{3F}\Big{[}(Fx_{1}-E_{0})^{3/2}-(Fa-E_{0})^{3/2}\Big{]}.$
Thus the exterior WKB tail behaves like $|u(x_{1})|\propto\exp(-d(a\to x_{1}))$, and since $d(a\to x_{1})\sim\mathrm{const}/F$ as $F\downarrow 0$, the leakage across the barrier —and hence the resonance width— is exponentially small in $1/F$. This explicit formula clarifies how the exponential suppression arises from the tunneling barrier created by the linear potential. In particular, it explains why the resonance width behaves like $\exp(-c/F)$ as $F\downarrow 0$. More precisely:
Theorem 15 (Exponentially small resonance widths for small fields).
Let $d=3$ and let $E_{0}<0$ be a simple eigenvalue of $H_{0,\alpha}$. There exist constants $F_{0},c_{1},c_{2}>0$ such that for all $0<F<F_{0}$ the operator $H_{F,\alpha}$ has a resonance $z(F)$ with $\text{{\rm Re}}z(F)\to E_{0}$ as $F\downarrow 0$ and
$e^{-c_{2}/F}\ \lesssim\ \Gamma(F):=-2\,\text{{\rm Im}}~z(F)\ \lesssim\ e^{-c_{1}/F}.$
The constants $c_{1},c_{2}$ can be expressed in terms of an Agmon-type distance along the field direction associated with the effective potential barrier.
Proof.
Let $E_{0}<0$ be a simple eigenvalue of $H_{0,\alpha}$ with normalized boundary eigenvector $\phi_{0}^{\flat}\in L^{2}(S_{a})$ as in Step 1. As explained above, the branch $z(F)$ of resonances near $E_{0}$ is obtained from the equation $G(F,z)=I+A_{\alpha}M_{F}(z)$ not invertible. Analytic perturbation theory yields
$z(F)=E_{0}+Fa_{1}+F^{2}a_{2}+O(F^{3}),\qquad a_{1}=-(M_{1}^{\flat}(E_{0})\phi_{0}^{\flat},\phi_{0}^{\flat})_{L^{2}(S_{a})},$
and $\text{{\rm Re}}z(F)\to E_{0}$ as $F\downarrow 0$. The imaginary part is more delicate.
By Herbst’s complex distortion, resonances of $H_{F,\alpha}$ correspond to discrete eigenvalues of $H_{F,\alpha}(\theta)$ for $0<\text{{\rm Im}}~\theta<\theta_{0}$. The corresponding eigenfunction $u_{F}$ satisfies $(H_{F,0}+A_{\alpha}\delta(|x|-a)-z(F))u_{F}=0$. Inside the wall, $u_{F}$ remains close to the bound state of $H_{0,\alpha}$, while in the exterior region $\{x_{1}>a\}$ it decays according to the Agmon action computed above. Applying Agmon’s method then yields the upper bound
$\Gamma(F)=-2\text{{\rm Im}}~z(F)\ \lesssim\ \exp(-c_{1}/F)$
for some $c_{1}>0$. Conversely, Carleman–type lower bounds exclude faster-than-exponential decay, giving
$\Gamma(F)\ \gtrsim\ \exp(-c_{2}/F)$
for some $c_{2}>0$. Both constants $c_{1},c_{2}$ are expressed in terms of the Agmon distance across the barrier region in the field direction.
Collecting these estimates, we obtain
$e^{-c_{2}/F}\ \lesssim\ \Gamma(F)\ \lesssim\ e^{-c_{1}/F},\qquad 0<F<F_{0},$
for some $F_{0}>0$. This proves the claim. $\Box$
This proof is a concise summary of the detailed arguments given by Wang [20, 21].
This establishes that bound states at $F=0$ bifurcate into resonances whose widths vanish faster than any power of $F$ as $F\downarrow 0$, in accordance with the general theory of metastable states under static fields.
Remark 5 (Physical interpretation for constant $\alpha<0$).
In the case of a constant attractive wall strength $\alpha<0$, the negative eigenvalue $E_{0}=-\kappa^{2}$ at $F=0$ corresponds to a bound state localized near the spherical shell. Switching on a small electric field $F>0$ tilts the potential and creates an effective barrier in the field direction. The Agmon action $d(a\to x_{1})$ measures the tunneling probability through this barrier, and the resonance width $\Gamma(F)$ corresponds to the decay rate of the metastable state. Thus the mathematical result that $\Gamma(F)\sim e^{-c/F}$ admits a direct physical interpretation: the bound state leaks out by quantum tunneling, in close analogy with the mechanism of $\alpha$–decay in nuclear physics.
Comparison with the free Stark Hamiltonian
It is instructive to compare our results with the well–known case of the free Stark Hamiltonian.
  • For the free Stark Hamiltonian, the spectrum is purely absolutely continuous and no bound states are present. Its resonances are known to exist only in an abstract sense, with widths that are exponentially small in the inverse field strength, but no explicit correspondence with eigenvalues is available.
  • By contrast, for the penetrable–wall Stark Hamiltonian studied here, bound states exist at zero field, and their continuation under a nonzero field yields resonances that can be traced explicitly. Their positions admit perturbative expansions, and their widths are again exponentially small, now governed by an Agmon–type distance determined by the wall and the associated bound state.
This comparison highlights that, unlike the free Stark Hamiltonian where resonances remain abstract, the penetrable–wall model provides a concrete and perturbative description of resonance formation and decay.
Point spectrum at zero field for constant wall strength
When the wall strength is constant, $\alpha(\omega)\equiv\alpha$, the negative spectrum of $H_{0,\alpha}$ can be determined explicitly in terms of modified spherical Bessel functions. For $l=0$, bound states at energy $E=-\kappa^{2}$ with $\kappa>0$ satisfy
$\alpha=-\kappa\bigl{(}1+\coth(\kappa a)\bigr{)}.$
From this relation one sees that only finitely many negative eigenvalues exist, accumulating at zero, in agreement with Ikebe–Shimada [14]. Note also that bound states appear only for attractive wall strength $\alpha<0$; for $\alpha>0$ the spectrum is purely $0,\infty)$.
Moreover, the equation yields simple asymptotics for $\kappa$:
  • In the weak coupling regime ($\kappa a\ll 1$),
    $\kappa\approx-\alpha-\frac{1}{a},\qquad E\approx-\left(\alpha+\frac{1}{a}\right)^{2}.$
  • In the strong coupling regime ($\kappa a\gg 1$),
    $\kappa\approx-\frac{\alpha}{2},\qquad E\approx-\frac{\alpha^{2}}{4}.$
These explicit formulas clarify the structure of the point spectrum at zero field and provide the starting point for the resonance asymptotics under a nonzero Stark field.

### v2

@@ 5.5 Small–field asymptotics in three dimensions
The detailed calculation in the preceding subsection can be placed in a general framework. Let $E_{0}<0$ be a simple eigenvalue of $H_{0,\alpha}$ with normalized boundary eigenvector $\phi_{0}^{\flat}\in L^{2}(S_{a})$ satisfying
$\left((M_{0}^{\flat})^{\prime}(E_{0})\phi_{0}^{\flat},\phi_{0}^{\flat}\right)_{L^{2}(S_{a})}=1.$
Then the implicit function theorem applied to the relation $I+A_{\alpha}M_{F}(z)$ gives an analytic branch
$z(F)=E_{0}+Fa_{1}+F^{2}a_{2}+O(F^{3}).$
In the ss–wave case we have already seen that $a_{1}=0$ and that $a_{2}$ is given by an explicit formula. In the constant–wall, ss–wave case one may take the normalized boundary eigenvector as $\phi_{0}^{\flat}=|S_{a}|^{-1/2}\chi_{S_{a}}$ (i.e. $(4\pi a^{2})^{-1/2}\chi_{S_{a}}$ in $d=3$), so that the abstract formula reduces exactly to the expression for $a_{2}$ obtained in Section 5.4 (see the display for $a_{2}$ before the Agmon action). Thus the general framework here reproduces the concrete computation in the ss–wave example.
The exponential smallness of the width follows from the Agmon distance in the field direction. In the exterior region one computes
$d(a\to x_{1})=\int_{a}^{x_{1}}\sqrt{Fs-E_{0}}\,ds,$
which behaves like $\tfrac{2}{3F}|E_{0}|^{3/2}$ as $F\downarrow 0$. Thus the resonance width satisfies
$e^{-c_{2}/F}\ \lesssim\ \Gamma(F)\ \lesssim\ e^{-c_{1}/F},$
for suitable $c_{1},c_{2}>0$. In the constant wall case these constants agree with $\tfrac{4}{3}|E_{0}|^{3/2}$, consistent with the explicit evaluation in the previous subsection.
It is instructive to compare this result with the free Stark Hamiltonian. In the free case the spectrum is purely absolutely continuous and no bound states exist. Resonances can only be defined abstractly and have widths that are exponentially small in $1/F$, but no explicit continuation of eigenvalues is available. In contrast, in the penetrable–wall case bound states at $F=0$ continue to resonances for $F\neq 0$. Their positions admit perturbative expansions and their widths are governed by the barrier action. This shows how localization due to the wall competes with delocalization in the field.
Appendix: Supplementary computations
This appendix contains technical details which support the results in Sections 5.4 and 5.5. In particular we give the representation of the matrix elements that appear in the quadratic shift and we outline the derivation of the prefactor in the resonance width.
A.1. Matrix elements for the quadratic shift (detailed)
We compute $\langle Y_{00},M_{1}(E_{0})Y_{10}\rangle$ and $\langle Y_{00},M_{2}(E_{0})Y_{00}\rangle$ in the constant wall case. By definition,
$M_{1}(z)=-\,\tau R_{0}(z)\,x_{1}\,R_{0}(z)\tau^{*},\qquad M_{2}(z)=\tau R_{0}(z)\,x_{1}R_{0}(z)\,x_{1}R_{0}(z)\tau^{*},$
and the free resolvent kernel at $z=-\kappa^{2}$ is
$G_{0}(x,y;z)=\frac{e^{-\kappa|x-y|}}{4\pi|x-y|}.$
Hence the boundary kernels on $S_{a}\times S_{a}$ are given by
$M_{1}(a\omega,a\omega^{\prime};z)$ $=-\int_{\mathbb{R}^{3}}G_{0}(a\omega,x;z)\,x_{1}\,G_{0}(x,a\omega^{\prime};z)\,dx,$
$M_{2}(a\omega,a\omega^{\prime};z)$ $=\int_{\mathbb{R}^{3}}G_{0}(a\omega,x;z)\,x_{1}\,G_{0}(x,y;z)\,x_{1}\,G_{0}(y,a\omega^{\prime};z)\,dx\,dy.$
We expand $G_{0}$ in spherical harmonics via the addition theorem on spheres of radius $r$ and $a$,
$G_{0}(a\omega,r\Omega;z)=\frac{\kappa}{4\pi}\sum_{\ell=0}^{\infty}(2\ell+1)\,i_{\ell}(\kappa r_{<})\,k_{\ell}(\kappa r_{>})\,P_{\ell}(\omega\cdot\Omega),$
and similarly for $G_{0}(r\Omega,a\omega^{\prime};z)$, where $r_{<}=\min\{r,a\}$ and $r_{>}=\max\{r,a\}$. We also write
$x_{1}=r\cos\theta=r\sqrt{\frac{4\pi}{3}}\,Y_{10}(\Omega).$
The angular integrals over $\Omega\in S^{2}$ reduce by the Gaunt coefficients
$\int_{S^{2}}Y_{\ell m}(\Omega)\,Y_{10}(\Omega)\,Y_{\ell^{\prime}m^{\prime}}(\Omega)\,d\Omega$
$=$ $\sqrt{\frac{(2\ell+1)3(2\ell^{\prime}+1)}{4\pi}}\,\bigl{\langle}\ell,0;1,0\,\big{|}\,\ell^{\prime},0\bigr{\rangle}\bigl{\langle}\ell,m;1,0\,\big{|}\,\ell^{\prime},m^{\prime}\bigr{\rangle},$
so that only the channels $\ell^{\prime}=\ell\pm 1$ contribute. Contracting with $Y_{00}(\omega)$ and $Y_{10}(\omega^{\prime})$ leaves the single channel $\ell=0$, $\ell^{\prime}=1$. After the $\Omega$–integration one obtains
$\langle Y_{00},M_{1}(E_{0})Y_{10}\rangle=-\frac{\kappa_{0}^{2}}{3}\,\frac{a^{2}}{(4\pi)}\int_{0}^{\infty}r^{3}\,\Xi_{1}(\kappa_{0};r,a)\,dr,$
where
$\Xi_{1}(\kappa;r,a)=\bigl{[}i_{0}(\kappa r_{<})k_{0}(\kappa r_{>})\bigr{]}\bigl{[}i_{1}(\kappa r_{<})k_{1}(\kappa r_{>})\bigr{]}.$
Splitting the radial integral at $r=a$ and using the recurrence and Wronskian identities for modified spherical Bessel functions,
$\frac{d}{dt}\bigl{(}t\,i_{\ell}(t)\bigr{)}$ $=t\,i_{\ell-1}(t)-\ell\,i_{\ell}(t),$
$\frac{d}{dt}\bigl{(}t\,k_{\ell}(t)\bigr{)}$ $=-t\,k_{\ell-1}(t)-\ell\,k_{\ell}(t),$
$i_{\ell}^{\prime}(t)k_{\ell}(t)-i_{\ell}(t)k_{\ell}^{\prime}(t)$ $=-\frac{1}{t^{2}}.$
one can integrate by parts to evaluate the radial integral in terms of the boundary value at $r=a$. This yields the compact form
$\langle Y_{00},M_{1}(E_{0})Y_{10}\rangle=-\frac{a}{3}\,Q_{1}(\kappa_{0}a),$
with
$Q_{1}(t)=t\bigl{(}i_{0}(t)k_{1}(t)+i_{1}(t)k_{0}(t)\bigr{)}.$
In the same manner, for $M_{2}$ we expand both $x_{1}$ factors and use the same selection rules. After the angular reduction and the same type of radial calculation, one obtains
$\langle Y_{00},M_{2}(E_{0})Y_{00}\rangle=\frac{a^{2}}{3}\,Q_{2}(\kappa_{0}a),$
where
$Q_{2}(t)=i_{0}(t)k_{0}(t)+t\bigl{(}i_{0}(t)k_{1}(t)+i_{1}(t)k_{0}(t)\bigr{)}.$
These formulas agree with the structure used in Section 5.4 and lead directly to the expression of $a_{2}$ in (5.3) after inserting the normalization on $L^{2}(S_{a})$ and the relation for $\mu_{0}^{\prime}(E_{0})$ derived there.
A.2. Prefactor in the resonance width
The exponential behavior of the resonance width is determined by the Agmon action $S(F)$, but the prefactor requires a more delicate argument. In the exterior region $\{x_{1}>a\}$ the radial ss–wave function satisfies
$u^{\prime\prime}(x_{1})+Fx_{1}u(x_{1})=Eu(x_{1}),\qquad E\approx E_{0}.$
By the scaling $y=F^{1/3}(x_{1}-E/F)$ this reduces to the Airy equation. The decaying solution is proportional to ${\rm Ai}(y)$, which controls the asymptotic tail of the resonance state.
Near the turning point $x_{1}=-E_{0}/F$ one applies the WKB connection formula, which connects the oscillatory solution inside the barrier with the decaying Airy tail outside. This matching determines the amplitude of the outgoing component at infinity. As a result one obtains
$\Gamma(F)\sim C(F)\exp(-2S(F)),$
where $S(F)$ is given by (5.5) and $C(F)$ varies slowly with $F$. In particular $C(F)\to C_{0}>0$ as $F\downarrow 0$. Thus the essential exponential factor is $\exp(-2S(F))$, and the prefactor only changes the constant in front.
This analysis confirms the expression of the resonance width in Section 5.4 and clarifies the role of the WKB connection at $r=a$ for the $Y_{00}$ component.